OpenStudy (anonymous):

ln(x^4) ln(x^4) @Mathematics

OpenStudy (anonymous):

Anyone out there?

OpenStudy (asnaseer):

what is your question?

OpenStudy (slaaibak):

4ln(x) ?

OpenStudy (chaise):

Slaabaik is right.

OpenStudy (anonymous):

ln(x^4+y^4) differentiate wrt x and y seperately

OpenStudy (slaaibak):

(4x^3 + dy/dx 4y^3)/(x^4 + y^4)

OpenStudy (anonymous):

differentiate wrt x only; i will figure the other out pls

OpenStudy (slaaibak):

must I differentiate ln(x^4)? okay. answer is: 4x^3/x^4 = 4/x

OpenStudy (anonymous):

@slaaibak no

OpenStudy (anonymous):

f=ln(x^4+y^4) differentiate wrt x

OpenStudy (slaaibak):

4x^3/(y^4 + x^4)

OpenStudy (asnaseer):

\[z=\ln (x^{4}+y^{4})\]this can be done in parts, so let:\[w=x^{4}\]then:\[z=\ln(w+y^{4})\]\[dz/dx=dz/dw*dw/dx=\frac{1}{w+y^{4}}*4x^{3}=\frac{4x^{3}}{x^{4}+y^{4}}\]

OpenStudy (anonymous):

Thanks guyz